OpenStudy (supercait10):

Question 1. Which statement is an example of the Identity Example Property of Multiplication? A. 8.0= 0 B. 8.1=8 C. 8.-1=-8 D. -8.1=8 Question 2. what is another way to write -3.(4+7)? A. -3.4+7 B. -3.4.7 C. -4.4+4.7 D. -3. 4+(-3).7 Question 3. Which property is represented by 5+(-8)=8+5? A. Identify B. Associative C. Commutative D. Distributive Question 4. What is the solution to any problem that involves the Multiplication Property of Zero? A.1 B.0 C. -1 D. -2 How could your rewrite this expression (-3+5) + 8+(-1) using the Associative Property of Addition?

OpenStudy (supercait10):

A. -3-8+(5+1) B. (-3+(-1)) + (5+8) C. -3+(5+8)+(-1) D. 5+8+3+1 For question Five

OpenStudy (anonymous):

this is pre algebra or ??

OpenStudy (supercait10):

yup pre algebra

OpenStudy (anonymous):

okay, I give you #4 is B :) now #1, D is wrong :) and Identify example of multiplication is a times 1 = a :) not 0 nor -1 , so what do you think?

OpenStudy (anonymous):

now look at #2 we have distribute multiplication :) so....... is will be D since we are multiplying -3 to both 4 and 7 not just one of them :)

OpenStudy (supercait10):

okay

OpenStudy (anonymous):

#3, the answer is B, why? cuz it give you the same answers, if you don't understand, read the definition of it. Lastly, #5 since I have explain above so basically that you just calculate the results of the given equations then calculate all equations from the giving results, if it match to the results that you have in the given equations, then it'll be your answer

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

well it looks like your covered

OpenStudy (anonymous):

do you understand? or you need me to show you steps?

OpenStudy (supercait10):

steps please

OpenStudy (anonymous):

for what# ?

OpenStudy (supercait10):

5

OpenStudy (anonymous):

okay, so you will have : (-3+5) + 8+(-1) = 2 + 7 = 9 ( I think you know how i got the results, if no, please let me know) A. -3-8+(5+1) = -11 + 6 = -5 ( false ) B. (-3+(-1)) + (5+8) = -4 + 13 = 9 ( true ) C. -3+(5+8)+(-1) = -3 + 13 + (-1) = 10 + (-1)= 9 ( true) D. 5+8+3+1= 17 so now you know that you have 2 correct results, what will you do? of course is chose one of them

OpenStudy (supercait10):

umm b

OpenStudy (anonymous):

but which one is right, I dun know to since your test give 2 correct answers :)

OpenStudy (supercait10):

ehh oh well i'll just message my teacher saying Question five has two correct answers

OpenStudy (anonymous):

yeah :3 do that :3 but show your work, okay?

OpenStudy (supercait10):

k